An elevator whose floor to ceiling height is 2.75m starts ascending with a constant upward acceleration of 1.2 m/s2. Two seconds after the start, a loose bolt drops from the ceiling towards the floor of the elevator. Calculate:

a)The time taken by the bolt to reach the floor

b)Displacement of the bolt during that time

c)The distance travelled by the bolt during the same time


1
Expert's answer
2021-08-11T07:28:42-0400

Part a

Relative velocity is zero

"s= ut +\frac{1}{2}at^2\\
2.75= 0*t +\frac{1}{2}*11*t^2\\
\implies t = 0.7 s"


Part b

"\Delta y = V_0 t+ \frac{1}{2}gt^2\\
\Delta y = 2.4*0.7+ \frac{1}{2}*9.81*0.7^2\\
\Delta y = -0.7 m"


Part c

"v^2=u^2-2gs\\
s= \frac{v^2}{2g}\\
s= \frac{2.4^2}{2*9.81}\\
s= 0.29 m\\
Total = 2*0.29 +0.7 \\
Total = 1.3 m"